Write the vector v=( 1, -2, 5) as a linear combination of the vector e1=( 1, 1, 1), e2= ( 1, 2, 3) and e3=( 2, -1, 1)
1
Expert's answer
2020-11-11T19:09:23-0500

linear combimation has the form:

"v \u20d7=\u03b1*(e1) \u20d7+\u03b2*(e2) \u20d7+\u03b3*(e3) \u20d7"


"(1; -2; 5)=\u03b1*(1;1;1)+\u03b2*(1;2;3)+\u03b3*(2; -1;1)"

"(1; -2; 5)=(\u03b1;\u03b1;\u03b1;)+(\u03b2;2\u03b2;3\u03b2)+(2\u03b3;-\u03b3;\u03b3)"

find "\alpha, \beta, \gamma:"

system of equations:


"\u03b1+\u03b2+2\u03b3=1, (1)""\u03b1+2\u03b2-\u03b3=-2, (2)""\u03b1+3\u03b2+\u03b3=5; (3)"


solution system:

equation (1) minus equation (2):

"-\beta+3\gamma=3;"

"\beta=3\gamma-3;" (4)

equation (2) plus equation (3):

"2\alpha+5\beta=3;"

"\u03b1=(3-5\u03b2)\/2;"

"\alpha=(3-5(3\gamma-3))/2;" (5)

equation (1):

"(3−5(3γ−3))/2+3\gamma-3+2\gamma=1;"

"3-15\gamma+15+10\gamma-6=2;"

"-5\gamma=-10;"

"\gamma=2;"

equation (4):

"\beta=3*2-3;"

"\beta=3;"

equation (1):

"\alpha=1-2\gamma-\beta;"

"\alpha=1-2*2-3;"

"\alpha=-6;"


"v \u20d7=-6*(e1) \u20d7+3*(e2) \u20d7+2*(e3);"

"(1; -2; 5)=-6*(1;1;1)+3*(1;2;3)+2*(2; -1;1);"